Fifteen litres of mixture contain water and milk in the ratio 2:4.if 3 litres of this mixture are replaced by 3 litres of water,what would be the ratio of water to milk in the new mixture?
1
Expert's answer
2020-10-20T17:28:06-0400

If 3 litres are removed from the mixture, 12 litres remain ie (15-3=12)

The 12 litres of mixture contain water and milk in the ratio 2:4

The amount of water in the mixture is "2\/(2+4)*12=4 litres"

The amount of milk is "4\/(2+4)*12=8litres"

Adding 3 litres of water to the mixture, amount of water becomes "4litres +3litres=7litres"

The new ratio of water to milk becomes "7litres:8litres =7:8"